40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Apr 15 18:59:41 2011, tara, DailyProgress, FSS, Preparing to install TTFSS 
    Reply  Tue Apr 19 22:40:03 2011, tara, DailyProgress, FSS, Preparing to install TTFSS 
       Reply  Wed Apr 20 17:54:52 2011, tara, DailyProgress, FSS, Preparing to install TTFSS 000_0025.jpg000_0024.jpg000_0023.jpgfastout.png
          Reply  Sat Apr 23 11:40:09 2011, tara, DailyProgress, FSS, Preparing to install TTFSS TTFSS_cartoon.pdf
             Reply  Sun Apr 24 22:02:07 2011, tara, DailyProgress, FSS, Preparing to install TTFSS beat_2011_04_24.pngbeat_2011_04_24.png
                Reply  Mon Apr 25 17:53:20 2011, tara, DailyProgress, FSS, Preparing to install TTFSS beat_2011_04_25.png
Message ID: 585     Entry time: Wed Apr 20 17:54:52 2011     In reply to: 584     Reply to this: 587
Author: tara 
Type: DailyProgress 
Category: FSS 
Subject: Preparing to install TTFSS 

TTFSS is not working properly yet. The signal from Fast out is distorted, but it still can lock the laser. All three actuators are being tested over night. Fine tuning has yet to be done. Once the temperature is settled, I'll measure the beat signal.

 

       I tried to lock the laser with the new TTFSS. There are three actuators, SLOW (thermal control), FAST (acts on PZT), and PC ( acts onEOM).  First, I locked the laser with FAST actuator only. PD input and LO input were removed from the old FSS card and connected to the TTFSS servo box.Then, I scanned the laser by modulating the laser frequency via laser's FAST input to check the error signal. The signal measured at mixer out, FAST MON, RAMP IN OUT2 (see fig 0, for where the outputs are) were good, see fig1 and fig2, but the signal from FAST OUT(tnc) is distorted, see fig3.

      Another problem was about the lemo output, it appeared that

     1)when I measured the signal from FAST OUT(lemo), I did not see any signal at all, and

     2)when I observed the distorted signal from FAST OUT(tnc) and plugged in a lemo connector to lemo FAST OUT, the signal from tnc also disappeared. The connector on the cable is the one we used for FSS, it should be working properly, but I will double check it. There might be shorted/ bad connections in the cable. It probably is the problem from the lemo connector on the board. When I did the test on the box 6 months ago, I used only FAST OUT(tnc).

  

fastout.png

fig0: a part of the schematic from TTFSS servo. FAST OUT has two output, tnc and lemo connectors. The full schematic can be found here.

I measured the signal from FASTM, RAMP IN OUT2, FAST OUT (tnc), and FAST OUT(lemo)

 

000_0025.jpg

 fig1: signal from mixer mon

 

000_0024.jpg

fig2 signal from fast out2 (similar to that from fast out mon (different in amplitude), so I post only this picture)

 

000_0023.jpg

fig3: signal from fast out (which is fed to the laser)

ELOG V3.1.3-